Showing posts with label probability. Show all posts
Showing posts with label probability. Show all posts

Tuesday, September 5, 2023

16.4 - Axiomatic Approach to Probability

In the previous section, we saw Exhaustive events. In this section, we will see Axiomatic approach to probability.

• In our earlier classes, we have seen how probability can be calculated in simple cases.
• We calculated probability as a fraction. It was a proper fraction. That is., numerator will be less than the denominator. If a proper fraction is converted to decimal form, the value on the left side of the decimal point will be zero.
• In the denominator, we gave total number of possible outcomes. In the numerator, we gave the total number of favorable outcomes.
• In our present discussion, we will see three axioms developed by the Russian mathematician A.N. Kolmogorov. Those axioms will enable us to calculate probability in complex problems.
• Axiom means a statement or rule, which is self-evident. Such a statement/rule can be a useful basis for more advanced mathematical calculations.

Axiom 1: For any event A, P(A) ≥ 0.
• This can be explained in 5 steps:
(i) The probability any event A to occur will be greater than or equal to zero.
    ♦ If P(A) is zero, A will never occur.
    ♦ If P(A) is large, there is a higher probability for A to occur.
    ♦ If P(A) is small, there is a lower probability for A to occur.
(ii) We saw that, probability is always a proper fraction. So the largest possible value for P(A) is 1. When the probability is 1, the event will surely occur.
(iii) Also, the smallest possible value for P(A) is zero. When the probability is zero, the event will never occur.
(iv) Since P(A) is a proper fraction, “zero value” is obtained when the numerator is zero. This happens when there are no favorable outcomes.
(v) Since P(A) is a proper fraction, “1 value” is obtained when the numerator is same as the denominator. This happens when number of favorable outcomes is same as the total number of outcomes.

Axiom 2:
For any experiment, P(S) = 1
• This can be explained in 2 steps:
(i) We know that, any event A can be denoted as a set A. The elements of set A will be the favorable outcomes of event A.
(ii) The sample space of an experiment is also a set. We denote it as S.
• So this S can be considered as an event. Whenever we perform that experiment, we will surely obtain an element of S. Based on this, we can say: Probability for S to occur is 1.

Axiom 3:
If A and B are two mutually exclusive events then P(A∪B) = P(A) + P(B)
• This can be explained using an example. It can be written in 5 steps:
(i) First we will assign the sets.
• Let set S of an experiment have 6 elements.
• Let event A of that experiment have 3 elements.
• Let event B of that experiment have 2 elements.
(ii) Now we will write the probabilities of the individual events:
• We get:
    ♦ P(A) = 3/6 = 1/2
    ♦ P(B) = 2/6 = 1/3
(iii) Given that A and B are disjoint sets. So A∪B will have 5 elements.
• So P(A∪B) = 5/6
(iv) Let us add the individual probabilities:
P(A) + P(B) = 1/2 + 1/3 = 5/6
• This is same as the result in (iii)
(v) Comparing the results in (iii) and (iv), we can write:
P(A∪B) = P(A) + P(B)


Using the above axioms, we can derive four important results.
Result 1:
This can be written in 3 steps:
1. Suppose that, B is a null set. That means, no outcome in S can be used to define the event B.
(For example, when a die is thrown once, the event of “getting number 7” is a null set)
• So we can write: B = ะค
2. Applying axiom 3, we get:
$\begin{array}{ll}
{}&{P(A \cup B)}
& {~=~}& {P(A) + P(B)}
&{} \\

{\Rightarrow}&{P(A \cup \phi)}
& {~=~}& {P(A) + P(\phi)~~ \color {green} {\text{- - - (I)}}}
&{} \\

{\Rightarrow}&{P(A)}
& {~=~}& {P(A) + P(\phi)~~ \color {green} {\text{- - - (II)}}}
&{} \\

{\Rightarrow}&{P(\phi)}
& {~=~}& {0}
&{} \\

\end{array}$

◼ Remarks:
• Line marked as I:
In this line, we are able to apply axiom 3 because, any set A and the null set are disjoint sets
• Line marked as II:
In this line, we use the rule:
Union of any set A with the null set will give A
3. Based on the above calculations, we can write:
If there are no outcomes for an event, then the probability for that event to occur is zero.


Result 2:
This can be written in 4 steps:
1. Let S be the sample space of an experiment.
• Also, let ๐œ”1, ๐œ”2, ๐œ”3, . . . , ๐œ”n be the n elements of S.
• Then we can write: S= {๐œ”1, ๐œ”2, ๐œ”3, . . . , ๐œ”n}
2. So S is the union of n singleton sets:
S = {๐œ”1} ∪ {๐œ”2} ∪ {๐œ”3} ∪ . . . ∪ {๐œ”n}
3. Applying axiom 3, we get:

$\begin{array}{ll}
{}&{\rm{P \left(\lbrace \omega_1 \rbrace \cup \lbrace \omega_2 \rbrace \cup \lbrace \omega_3 \rbrace \cup ~.~.~.~ \cup \lbrace \omega_n \rbrace \right)}}
& {~=~}& {\rm{P \left(\lbrace \omega_1 \rbrace \right) + P \left(\lbrace \omega_2 \rbrace \right) + P \left(\lbrace \omega_3 \rbrace \right) ~+~.~.~.~+~P \left(\lbrace \omega_n \rbrace \right)}}
&{} \\

{\Rightarrow}&{\rm{P(S)}}
& {~=~}& {\rm{P \left(\lbrace \omega_1 \rbrace \right) + P \left(\lbrace \omega_2 \rbrace \right) + P \left(\lbrace \omega_3 \rbrace \right) ~+~.~.~.~+~P \left(\lbrace \omega_n \rbrace \right)}}
&{} \\

{\Rightarrow}&{\rm{1}}
& {~=~}& {\rm{P \left(\lbrace \omega_1 \rbrace \right) + P \left(\lbrace \omega_2 \rbrace \right) + P \left(\lbrace \omega_3 \rbrace \right) ~+~.~.~.~+~P \left(\lbrace \omega_n \rbrace \right)}}
&{} \\

\end{array}$

◼ Remarks:
From axiom 2, we know that: P(S) = 1

4. So we can write:
Sum of the individual probabilities of all individual outcomes of an experiment will be 1.


Result 3:
This can be written in 5 steps:
1. From result 2, we know that:
P({๐œ”1}) + P({๐œ”2}) + P({๐œ”3}) + . . . + P({๐œ”n}) = 1
• Note that, the sum on the RHS is 1.
2. Each term on the LHS is a proper fraction.
• Since the sum is 1, each term must be either less than 1 or equal to 1.
3. But axiom 1 tells us that:
Each term will be either greater than zero or equal to zero.
4. Combining (2) and (3), we get:
0 ≤ Any term ≤ 1
• Mathematically, we write this as:
0 ≤ P({๐œ”i}) ≤ 1
5. So we can write:
Take any outcome from S. That outcome can be considered as a simple event (event with only one outcome). The probability of that event will be greater than or equal to zero and at the same time, less than or equal to 1.


Result 4:
This can be written in 4 steps:
1. Let A be an event in an experiment.
• Also, let ๐œ”1, ๐œ”2, ๐œ”3, . . . , ๐œ”n be the n elements of A.
• Then we can write: A = {๐œ”1, ๐œ”2, ๐œ”3, . . . , ๐œ”n}
2. So A is the union of n singleton sets:
A = {๐œ”1} ∪ {๐œ”2} ∪ {๐œ”3} ∪ . . . ∪ {๐œ”n}
3. Applying axiom 3, we get:

$\begin{array}{ll}
{}&{\rm{P \left(\lbrace \omega_1 \rbrace \cup \lbrace \omega_2 \rbrace \cup \lbrace \omega_3 \rbrace \cup ~.~.~.~ \cup \lbrace \omega_n \rbrace \right)}}
& {~=~}& {\rm{P \left(\lbrace \omega_1 \rbrace \right) + P \left(\lbrace \omega_2 \rbrace \right) + P \left(\lbrace \omega_3 \rbrace \right) ~+~.~.~.~+~P \left(\lbrace \omega_n \rbrace \right)}}
&{} \\

{\Rightarrow}&{\rm{P(A)}}
& {~=~}& {\rm{P \left(\lbrace \omega_1 \rbrace \right) + P \left(\lbrace \omega_2 \rbrace \right) + P \left(\lbrace \omega_3 \rbrace \right) ~+~.~.~.~+~P \left(\lbrace \omega_n \rbrace \right)}}
&{} \\

{\Rightarrow}&{\rm{P(A)}}
& {~=~}& {\rm{\sum{P \left(\lbrace \omega_i \rbrace \right)}}}
&{} \\

\end{array}$

4. So we can write:
Sum of the individual probabilities of all individual outcomes of an event will be the probability of that event.


◼ Based on the three axioms and four results, we can write:
Any experiment must satisfy two conditions:
Condition 1:
Take any individual outcome of an experiment. The probability of that outcome will be greater than or equal to zero, but at the same time less than or equal to 1.

• Mathematically, we can write this as:
$\rm{0 \le P(\lbrace \omega_i\rbrace) \le 1,~Where~ \omega_i \in S}$
Condition 2:
Take the sum of the probabilities of all individual outcomes in S. That sum will be 1.

• Mathematically, we can write this as:
$\rm{\sum{P(\lbrace \omega_i\rbrace)} = 1,~Where~ \omega_i \in S}$

• Let us see an example. It can be written in steps:
1. Consider the experiment of tossing a coin once.
We know that, S = {H,T}
2. Let us assign individual probabilities for the two outcomes:
    ♦ P({H}) = ½
    ♦ P({T}) = ½
3. Does the above probability values satisfy the two conditions?
(i) We have: 0 ≤ ½ ≤ 1
So condition 1 is satisfied.
(ii) We have: (½ + ½) = 1
So condition 2 is also satisfied.
4. We know that exact ½ cannot be obtained even if we repeat the experiment 50000 times. What if we repeat the experiment only 20 times and get the following values?
    ♦ P({H}) = 1/4
    ♦ P({T}) = 3/4
5. Does the above probability values satisfy the two conditions?
(i) We have:
    ♦ 0 ≤ 1/4 ≤ 1
    ♦ 0 ≤ 3/4 ≤ 1
So condition 1 is satisfied.
(ii) We have: (1/4 + 3/4) = 1
So condition 2 is also satisfied.
6. We see that ¼ and ¾ also satisfy the two conditions.
• In fact, there are infinite number of possible values such that:
    ♦ P({H}) = p
    ♦ P({T}) = (1- p)
Where 0 ≤ p ≤ 1


Now we will see a solved example:

Solved example 16.9
Let a sample space be S = {๐œ”1, ๐œ”2, ๐œ”3, . . . , ๐œ”6}. Which of the following assignments of probabilities to each outcome are valid ?

$\begin{array}{cc}
{\textbf{Outcomes}}&{\mathbf{\omega_1}}
& {\mathbf{\omega_2}}& {\mathbf{\omega_3}}
&{\mathbf{\omega_4}} &{\mathbf{\omega_5}} &{\mathbf{\omega_6}}\\

{\textbf{(a)}}&{\frac{1}{6}}
& {\frac{1}{6}}& {\frac{1}{6}}
&{\frac{1}{6}} &{\frac{1}{6}} &{\frac{1}{6}}\\

{\textbf{(b)}}&{1}
& {0}& {0}
&{0} &{0} &{0}\\

{\textbf{(c)}}&{\frac{1}{8}}
& {\frac{2}{3}}& {\frac{1}{3}}
&{\frac{1}{3}} &{- \frac{1}{4}} &{- \frac{1}{3}}\\

{\textbf{(d)}}&{\frac{1}{12}}
& {\frac{1}{12}}& {\frac{1}{6}}
&{\frac{1}{6}} &{\frac{1}{6}} &{\frac{3}{2}}\\

{\textbf{(e)}}&{0.1}
& {0.2}& {0.3}
&{0.4} &{0.5} &{0.6}\\

\end{array}$

Solution:
We have two conditions to check whether the probabilities assigned to the outcomes are valid.
Condition 1:
$\rm{0 \le P(\lbrace \omega_i\rbrace) \le 1,~Where~ \omega_i \in S}$
Condition 2:
$\rm{\sum{P(\lbrace \omega_i\rbrace)} = 1,~Where~ \omega_i \in S}$

Let us check each case:
Part (a):
• Condition 1 is satisfied because, all probability values are greater than or equal to 0 and at the same time, less than or equal to 1.
• Condition 2 is satisfied because:
$\frac{1}{6} + \frac{1}{6} + \frac{1}{6} + \frac{1}{6} + \frac{1}{6} + \frac{1}{6}~=~1$
• Both conditions are satisfied. So the assigned probabilities are valid.

Part (b):
• Condition 1 is satisfied because, all probability values are greater than or equal to 0 and at the same time, less than or equal to 1.
• Condition 2 is satisfied because:
1 + 0 + 0 + 0 + 0 + 0 = 1
• Both conditions are satisfied. So the assigned probabilities are valid.

Part (c):
• Condition 1 is not satisfied because, P({๐œ”5}) and P({๐œ”6}) are less than 0.
• Since condition 1 is not satisfied, there is no need to check condition 2.
• Condition 1 is not satisfied. So the assigned probabilities are not valid.

Part (d):
• Condition 1 is not satisfied because, P({๐œ”6}) is greater than 1.
• Since condition 1 is not satisfied, there is no need to check condition 2.
• Condition 1 is not satisfied. So the assigned probabilities are not valid.

Part (e):
• Condition 1 is satisfied because, all probability values are greater than or equal to 0 and at the same time, less than or equal to 1.
• Condition 2 is not satisfied because:
0.1 + 0.2 + 0.3 + 0.4 + 0.5 + 0.6 = 2.1 ≠ 1
• Condition 2 is not satisfied. So the assigned probabilities are not valid.


In the next section, we will see Probability of an event.

Previous

Contents

Next

Copyright©2023 Higher secondary mathematics.blogspot.com

Tuesday, August 29, 2023

16.3 - Exhaustive Events

In the previous section, we saw algebra of events. We also saw mutually exclusive events. In this section, we will see Exhaustive events.

Exhaustive events

This can be explained in 7 steps:
1. Consider the experiment of rolling a die.
• We know that S = {1, 2, 3, 4, 5, 6}
2. Suppose that event A occurs if the number obtained is less than 4.
• Then we have: A = {1, 2,3}
3. Suppose that event B occurs if the number obtained is greater than 2 but less than 5.
• Then we have: B = {3,4}
4. Suppose that event C occurs if the number obtained is greater than 4.
• Then we have: C = {5,6}
5. Let us find A∪B∪C:
{1,2,3}∪{3,4}∪{5,6} = {1,2,3,4,5,6} = S
• We see that:
The union is same as S
6. Now we can write the definition of exhaustive events. It can be written in 2 steps:
(i) Let E1, E2, E3, . . . En be some events associated with an experiment.
(ii) If the union of those events give the S of that experiment, then those events are called exhaustive events.
7. Let us see an interesting fact about exhaustive events. It can be written in 3 steps:
(i) Let E1, E2, E3, . . . En be the exhaustive events associated with an experiment.
• The union of those events will be S. We can write:
$\rm{E_1 \cup E_2 \cup E_3 \cup~.~.~.~ \cup E_n~=~\cup_{i=1}^{i=n}{E_i}~=~S}$
(ii) We know that:
• When a set is formed by the union of two or more sets,
    ♦ All elements of all participating sets
    ♦ Will be present in the resulting set.
• So all elements of E1, E2, E3, . . . En , will be present in $\rm{\cup_{i=1}^{i=n}{E_i}}$.
(iii) We also know that:
Whenever we perform the experiment, the outcome will be a member of S.
• But $\rm{S\,=\,\cup_{i=1}^{i=n}{E_i}}$
• So we can write:
Whenever the experiment is performed, one of the exhaustive events will surely occur.


Events which are both mutually exclusive and exhaustive

This can be explained in 3 steps:
1. Let E1, E2, E3, . . . En be some exhaustive events associated with an experiment.
2. Take any two sets from that list of exhaustive events. Those two sets must be disjoint.
• That is., which ever pair we take, the sets in that pair must be disjoint.
3. If the condition in (2) is satisfied, then we can say:
E1, E2, E3, . . . En are both mutually exclusive and exhaustive.


Now we will see some solved examples
Solved example 16.7
Two dice are thrown and the sum of the numbers which come up on the dice is noted. Let us consider the following events associated with the experiment.
A: the sum is even
B: the sum is a multiple of 3
C: the sum is less than 4
D: the sum is greater than 11
Which pairs of these events are mutually exclusive?
Solution:
1. We know that, the sample space is:
S ={
(1,1), (1,2), (1,3), (1,4), (1,5), (1,6),
(2,1), (2,2), (2,3), (2,4), (2,5), (2,6),
(3,1), (3,2), (3,3), (3,4), (3,5), (3,6),
(4,1), (4,2), (4,3), (4,4), (4,5), (4,6),
(5,1), (5,2), (5,3), (5,4), (5,5), (5,6),
(6,1), (6,2), (6,3), (6,4), (6,5), (6,6)}
2. Now we can write the events:
(i) A: the sum is even
So A ={
(1,1), (1,3), (1,5),
(2,2), (2,4), (2,6),
(3,1), (3,3), (3,5),
(4,2), (4,4), (4,6),
(5,1), (5,3), (5,5),
(6,2), (6,4), (6,6)}
(ii) B: the sum is a multiple of 3
So B ={
(1,2), (1,5),
(2,1), (2,4),
(3,3), (3,6),
(4,2), (4,5),
(5,1), (5,4),
(6,3), (6,6)}
(iii) C: the sum is less than 4
So C ={
(1,1), (1,2), (2,1)}
(iv) D: the sum is greater than 11
So D = {(6,6)}
3. We have four sets: A, B, C and D
We must consider all possible pairs.
(i) Consider A and B
They are not disjoint sets because, there are some common elements. For example, (1,5).
(ii) Consider A and C
They are not disjoint sets because, there is one common element which is (1,1).
(iii) Consider A and D
They are not disjoint sets because, there is one common element which is (6,6).
(iv) Consider B and C
They are not disjoint sets because, there are two common elements: (1,2) and (2,1).
(v) Consider B and D
They are not disjoint sets because, there is one common element which is (6,6).
(vi) Consider C and D
They are disjoint sets because, there are no common elements.
4. So we can write:
Only one pair "C, D" are mutually exclusive events.

Solved example 16.8
A coin is tossed three times, consider the following events.
A: No head appears
B: Exactly one head appears
C: Atleast two heads appear
Do they form a set of mutually exclusive and exhaustive events?
Solution:
1. We know that, the sample space is:
S = {(H,H,H), (H,H,T), (H,T,H), (H,T,T), (T,H,H), (T,H,T), (T,T,H), (T,T,T)}
2. Now we can write the events:
(i) A: No head appears
So A = {(T,T,T)}
(ii) B: Exactly one head appears
So B = {(H,T,T), (T,H,T), (T,T,H)}
(iii) C: Atleast two heads appear
So C = {(H,H,H), (H,H,T), (H,T,H), (T,H,H)}
3. First we will check whether they are exhaustive events. For that, we must find A∪B∪C. We get:
{(T,T,T)}∪{(H,T,T), (T,H,T), (T,T,H)}∪ {(H,H,H), (H,H,T), (H,T,H), (T,H,H)}
= {(T,T,T), (H,T,T), (T,H,T), (T,T,H), (H,H,H), (H,H,T), (H,T,H), (T,H,H)}
• Compare this union with
S = {(H,H,H), (H,H,T), (H,T,H), (H,T,T), (T,H,H), (T,H,T), (T,T,H), (T,T,T)}
• We see that, A∪B∪C = S
So A, B and C are exhaustive events.
4. Next we check whether A, B and C are mutually exclusive. For that, we must consider all possible pairs.
(i) Consider A and B
They are disjoint sets because, there are no common elements.
(ii) Consider A and C
They are disjoint sets because, there are no common elements.
(iii) Consider B and C
They are disjoint sets because, there are no common elements.
5. So we can write:
A, B and C are mutually exclusive and exhaustive events.


Link to a few more solved examples is given below:

Exercise 16.2


In the next section, we will see Axiomatic approach to probability.

Previous

Contents

Next

Copyright©2023 Higher secondary mathematics.blogspot.com

Saturday, August 26, 2023

Chapter 16.1 - Event in Probability

In the previous section, we completed a discussion on sample space. In this section, we will see event.

Event can be explained in 9 steps:
1. Consider the experiment of tossing a coin twice.
• We know that the sample space is:
S = {(H,H), (H,T), (T,H), (T,T)}
2. Suppose that, we are interested in those outcomes in which H occurs exactly once.
• Then we can pick out two outcomes: (H,T) and (T,H).
• We can form a set E using the two outcomes that we picked out:
E = {(H,T), (T,H)}
• We can write 3 points:
(i) We are interested in those outcomes in which H occurs exactly once.
• So we say that:
In the experiment, if H occurs exactly once, we have an event.
(ii) Two outcomes are favorable for the event.
• We write a set E in such a way that, those two outcomes are the only elements.
(iii) Then E will be a subset of S.
3. Suppose that, we are interested in those outcomes in which T occurs exactly two times.
• Then we can pick out one outcome: (T,T).
• We can form a set E using the outcome that we picked out:
E = {(T,T)}
• We can write 3 points:
(i) We are interested in those outcomes in which T occurs exactly two times.
• So we say that:
In the experiment, if T occurs exactly two times, we have an event.
(ii) One outcome is favorable for the event.
• We write a set E in such a way that, that one outcome is the only element.
(iii) Then E will be a subset of S.
4. Suppose that, we are interested in those outcomes in which T occurs at least once.
• Then we can pick out three outcomes: (H,T), (T,H) and (T,T).
• We can form a set E using the three outcomes that we picked out:
E = {(H,T), (T,H), (T,T)}
• We can write 3 points:
(i) We are interested in those outcomes in which T occurs at least once.
• So we say that:
In the experiment, if T occurs at least once, we have an event.
(ii) Three outcomes are favorable for the event.
• We write a set E in such a way that, those three outcomes are the only elements.
(iii) Then E will be a subset of S.
5. Suppose that, we are interested in those outcomes in which number of H is atmost 1.
("atmost 1" means, the number must not exceed 1. In other words, maximum allowed is 1)
• Then we can pick out three outcomes: (H,T), (T,H) and (T,T).
• We can form a set E using the three outcomes that we picked out:
E = {(H,T), (T,H), (T,T)}
• We can write 3 points:
(i) We are interested in those outcomes in which number of H is atmost 1.
• So we say that:
In the experiment, if H occurs atmost one time, we have an event.
(ii) Three outcomes are favorable for the event.
• We write a set E in such a way that, those three outcomes are the only elements.
(iii) Then E will be a subset of S.
6. Suppose that, we are interested in those outcomes in which second toss is not head.
• Then we can pick out two outcomes: (H,T) and (T,T).
• We can form a set E using the two outcomes that we picked out:
E = {(H,T), (T,T)}
• We can write 3 points:
(i) We are interested in those outcomes in which second toss is not H.
• So we say that:
In the experiment, if second toss is not H, we have an event.
(ii) Two outcomes are favorable for the event.
• We write a set E in such a way that, those two outcomes are the only elements.
(iii) Then E will be a subset of S.
7. Suppose that, we are interested in those outcomes in which number of T is atmost 2.
• Then we can pick out all the four outcomes: (H,H), (H,T), (T,H) and (T,T).
• We can form a set E using the four outcomes that we picked out:
E = {(H,H), (H,T), (T,H), (T,T)}
• We can write 3 points:
(i) We are interested in those outcomes in which number of T is atmost 2.
• So we say that:
In the experiment, if number of T is atmost 2, we have an event.
(ii) Four outcomes are favorable for the event.
• We write a set E in such a way that, those four outcomes are the only elements.
(iii) Then E will be a subset of S.
8. Suppose that, we are interested in those outcomes in which number of T is more than 2.
• Then we can pick out none of the four outcomes.
• We can form only a null set. A null set is also a subset of S.
• We can write 3 points:
(i) We are interested in those outcomes in which number of T is more than 2.
• So we say that:
In the experiment, if number of T is more than 2, we have an event.
(ii) No outcome is favorable for the event.
• We write a set E which is a null set.
(iii) Then E is a subset of S.
(Recall that, null set is also a subset)
9. Based on the above steps, we can write a definition for event. It can be written in 2 steps:
(i) An event is a set. It is denoted using the letter ‘E’
(ii) It is a subset of S.
• So all elements of E are outcomes.


Occurrence of an event

This can be explained in 4 steps:
1. Consider the experiment of rolling a die.
• We know that S = {1, 2, 3, 4, 5, 6}
2. We are interested in those outcomes which are less than 4.
• Then we can pick out three outcomes: 1, 2 and 3.
• We can write a set E using the three outcomes that we picked out:
E = {1, 2, 3}
3. Now we can write about the occurrence of the event:
• When the die is rolled, if 1 is obtained, then we say:
Event E has occurred.
• When the die is rolled, if 2 is obtained, then we say:
Event E has occurred.
• When the die is rolled, if 3 is obtained, then we say:
Event E has occurred.
4. Based on the above three steps, we can write the definition for “occurrence of event”. It can be written in 3 steps:
(i) Let an outcome ๐›š of an experiment occur.
(ii) Let ๐›š be an element of E. In other words, ๐›š ∈ E.
• Then we say that:
Event E has occurred.
(iii) If ๐›š ∉ E, then we say that:
Event E has not occurred.


Types of events

There are four types of events.
(i) Impossible event  (ii) Sure event  (iii) Simple event  (iv) Compound event.

Impossible event

This can be explained in 3 steps:
1. Consider the experiment of rolling a die.
• We know that S = {1, 2, 3, 4, 5, 6}
2. Suppose that, we are interested in those outcomes which are multiples of 7.
• Then we can pick out no outcomes.
• So the set E will be a null set.
3. If E is a null set, then that event is an impossible event.
• We can do the experiment any number of times we like. We will never get an outcome which is an element of E. Because, E is a null set.

Sure event

This can be explained in 3 steps:
 1. Consider the experiment of rolling a die.
• We know that S = {1, 2, 3, 4, 5, 6}
2. Suppose that, we are interested in those outcomes which are either odd or even.
• Then we can pick out all six outcomes.
• We can write a set E using those outcomes:
E = {1, 2, 3, 4, 5, 6}
   ♦ We see that, E is same as S.
3. If E = S, then that event is a sure event.
• We can do the experiment any number of times we like. The outcome will always be an element of E. Because, all outcomes are present in E.

Simple event

This can be explained in 7 steps:
1. Consider the experiment of tossing a coin two times.
We know that S = {(H,H), (H,T), (T,H), (T,T)}
2. Suppose that, we are interested in those outcomes in which:
   ♦ First toss gives T.
   ♦ Second toss gives H.
• Then we can pick out only one outcome, which is (T,H).
• We can write a set E using this outcome:
E = {(T,H)}
3. If E has only one element, then that event is called a simple event.
4. In fact, we can pick out each element from S and write distinct sets.
   ♦ E1 = {(H,H)}
   ♦ E2 = {(H,T)}
   ♦ E3 = {(T,T)}
• E1 is the event in which both tosses give H.
• E2 is the event in which first toss gives H and second toss gives T.
• E3 is the event in which both tosses give T.
5. So we can write an important point:
If there are n elements in S, then there will be n simple events.
6. Consider the event in which there is atleast one H.
• This event is not a simple event because, the set of this event has more than one elements.
7. A simple event is also known as an elementary event.    

Compound event

This can be explained in 5 steps:
1. Consider the experiment of tossing a coin three times.
We know that S = {(H,H,H), (H,H,T), (H,T,H), (H,T,T), (T,H,H), (T,H,T), (T,T,H), (T,T,T)}
(see first example in exercise 16.1 of the previous section)
2. Suppose that, we are interested in those outcomes in which:
   ♦ Exactly one H is obtained.
• Then we can pick out three outcomes, which are (H,T,T), (T,H,T) and (T,T,H).
• We can write a set E using these outcomes:
E = {(H,T,T), (T,H,T), (T,T,H)}
3. If E has more than one element, then that event is called a compound event.
4. We can write more examples from this experiment. Let us write two such examples:
Example (i):
Suppose that, we are interested in those outcomes in which:
   ♦ Atleast one H is obtained.
• Then we can write:
   ♦ E1 = {(H,H,H), (H,H,T), (H,T,H), (H,T,T), (T,H,H), (T,H,T), (T,T,H), (T,T,T)}
• E1 is a compound event because there are more than one elements.
Example (ii):
Suppose that, we are interested in those outcomes in which:
   ♦ Atmost one H is obtained.
• Then we can write:
   ♦ E2 = {(H,T,T), (T,H,T), (T,T,H), (T,T,T)}
• E2 is a compound event because there are more than one elements.
5. So we can write an important point:
All subsets of S, which have more than one elements, are compound events.


In the next section, we will see Algebra of Events.

Previous

Contents

Next

Copyright©2023 Higher secondary mathematics.blogspot.com

Monday, August 21, 2023

Chapter 16 - Probability

In the previous section, we completed a discussion on statistics. In this chapter, we will see probability.

In our earlier classes, we have seen the basics of probability. The links to those notes are given below:

    ♦ Probability part I consists of chapters 1.5, 1.6, . . . up to 1.9
    ♦ probability part II consists of chapters 28, 28.1, . . . up to 28.4
    ♦ probability part III consists of chapters 36, 37.1 and 36.2

The reader must have a thorough knowledge on the above three parts. In our present discussion, we will see some advanced techniques.

First we must become familiar with four basic terms:
(i) Random experiments, (ii) Outcomes  (iii) Sample space (iv) Sample point.

Random experiments

This can be explained in 4 steps:
1. Consider an experiment in which a student is analyzing the sum of interior angles of various triangles.
• He takes the first triangle and measure it’s three interior angles. He finds that the sum of those three angles is 180o   
• He takes the second triangle and measure it’s three interior angles. He finds that the sum of those three angles is 180o   
• He can take any number of triangles. The sum will be always 180o.
• So the experiment of “analyzing sum of interior angles of triangles” will always give a fixed result. We can predict the result. We can predict that, the sum will be 180o
2. Consider the experiment of tossing a coin.
• A student tosses a coin. The result can be head or tail. We cannot predict the result.
• The student tosses the coin again. This time also, the result can be head or tail. We cannot predict the result.
• In this experiment,
    ♦ There are two possible results: head and tail.
    ♦ It is not possible to predict the result.
• Such experiments are called random experiments.
• Note that, if there is only one possible result, a need for prediction will not arise.
• We can write:
An experiment is a random experiment if it satisfies two conditions:
(i) It has more than one possible result.
(ii) It is not possible to predict the result in advance.   
3. Based on the above steps, we can write:
The experiment of rolling a die is a random experiment.
• This is because, there are six possible results. Also, it is not possible to predict the result.
4. In this chapter, whenever we mention the word “experiment”, it will be a random experiment.

Outcomes  

This can be explained in 4 steps:
1. Consider the experiment of rolling a die.
2. The possible results are: 1, 2, 3, 4, 5 and 6
3. Each of the above possible results is called an outcome of the experiment.
4. So we can write:
    ♦ Outcomes of an experiment
    ♦ are
    ♦ possible results of that experiment.

Sample space

This can be explained in 4 steps:
1. Consider the experiment of rolling a die.
2. The outcomes are: 1, 2, 3, 4, 5 and 6
3. Using these outcomes, we can form a set S:
S = {1, 2, 3, 4, 5, 6}
• This set is called sample space of the experiment.
4. So we can define sample space in 3 steps:
(i) Sample space of an experiment, is a set
(ii) All outcomes of that experiment will be elements of that set.
(iii) This set is denoted by the letter ‘S’.

Sample point

This can be explained in 2 steps:
1. Each element of S is called a sample point.
2. But we know that, each element is an outcome.
• So we can write:
Each outcome is a sample point.


Let us see some solved examples:
Solved Example 16.1
Two different coins are tossed once. Find the sample space.
Solution:
1. Given that the coins are different.
    ♦ So the head of one coin will be different from the head of the other coin.
    ♦ Also, the tail of one coin will be different from the tail of the other coin.
• Because of this difference, we will give specific names:
    ♦ Head of first coin can be named as H1.
    ♦ Head of the other coin can be named as H2.
    ♦ Tail of first coin can be named as T1.
    ♦ Tail of the other coin can be named as H2.
2. Let us write the possible outcomes:
(i) Outcome 1:
    ♦ First coin gives H1.
    ♦ The other coin gives H2.
• Note that, in this outcome, T1 and T2 are not possible. For example, the first coin landed with H1 on the upper side. So T1 will not be visible.
• We can write this outcome as: (H1,H2)
(ii) Outcome 2:
    ♦ First coin gives H1.
    ♦ The other coin gives T2.
• Note that, in this outcome, T1 and H2 are not possible.
• We can write this outcome as: (H1,T2)
(iii) Outcome 3:
    ♦ First coin gives T1.
    ♦ The other coin gives H2.
• Note that, in this outcome, H1 and T2 are not possible.
• We can write this outcome as: (T1,H2)
(iv) Outcome 4:
    ♦ First coin gives T1.
    ♦ The other coin gives T2.
• Note that, in this outcome, H1 and H2 are not possible.
• We can write this outcome as: (T1,T2)
3. Based on the above outcomes, we can write the sample space as:
S = {(H1,H2), (H1,T2), (T1,H2), (T1,T2)}

Solved Example 16.2
Two different dice (one blue and the other red) are rolled once. Find the sample space.
Solution:
1. We have already seen this type of problem in our earlier classes (see solved example 1.14 in section 1.9).
2. In our present case, we are asked to write the sample space.
• For presenting the sample space in a “convenient and easy to understand” manner, we use ordered pairs (x,y)
• If we take the blue die as the first die, then:
    ♦ All values associated with blue die are x.
    ♦ All values associated with red die are y.
3. Based on such a notation, the sample space for our present case can be written as:
S ={
(1,1), (1,2), (1,3), (1,4), (1,5), (1,6),
(2,1), (2,2), (2,3), (2,4), (2,5), (2,6),
(3,1), (3,2), (3,3), (3,4), (3,5), (3,6),
(4,1), (4,2), (4,3), (4,4), (4,5), (4,6),
(5,1), (5,2), (5,3), (5,4), (5,5), (5,6),
(6,1), (6,2), (6,3), (6,4), (6,5), (6,6)}
• There are 36 elements in S.

Solved example 16.3
In each of the following experiments, write the sample space.
(i) A boy has a 1 rupee coin, a 2 rupee coin and a 5 rupee coin in his pocket. He takes two coins out of his pocket, one after the other.
(ii) A person is noting down the number of accidents along a busy highway during a year.
Solution:
Part (i):
1. We will use the form of ordered pairs (x,y)
    ♦ x is related to the first coin taken out.
    ♦ y is related to the second coin taken out.
2. Let us write the outcomes:
(i) First coin is 1, second coin is 2.
We write this as (1,2).
(ii) First coin is 1, second coin is 5.
We write this as (1,5)    
(iii) First coin is 2, second coin is 1.
We write this as (2,1)    
(iv) First coin is 2, second coin is 5.
We write this as (2,5)    
(v) First coin is 5, second coin is 1.
We write this as (5,1)    
(vi) First coin is 5, second coin is 2.
We write this as (5,2)
3. So the sample space can be written as:
S = {(1,2), (1,5), (2,1), (2,5), (5,1), (5,2)}
• There are 6 elements in S.
Part (ii):
1. The number of accidents in a year may be 0 (if there are no accidents), or any positive integer.
2. So the sample space can be written as:
S = {0, 1, 2, 3, . . . }

Solved example 16.4
A coin is tossed. If it shows head, we draw a ball from a bag containing 3 blue and 4 white balls. If it shows tail, we through a die. Write the sample space of this experiment.
Solution:
1. We will use the form of ordered pairs (x,y)
    ♦ x is related to tossing of coin.
    ♦ y is related to drawing a ball or rolling the die.
2. Let us write the out comes:
(i) The coin shows heads. The ball drawn is the first blue ball.
• We write this as (H,B1).    
(ii) The coin shows heads. The ball drawn is the second blue ball.
• We write this as (H,B2).    
(iii) The coin shows heads. The ball drawn is the third blue ball.
• We write this as (H,B3).
(iv) The coin shows heads. The ball drawn is the first white ball.
• We write this as (H,W1).
(v) The coin shows heads. The ball drawn is the second white ball.
• We write this as (H,W2).   
(vi) The coin shows heads. The ball drawn is the third white ball.
• We write this as (H,W3).    
(vii) The coin shows heads. The ball drawn is the fourth white ball.
• We write this as (H,W4).
(viii) The coin shows tails. The die shows 1.
We write this as (T,1).         
(ix) The coin shows tails. The die shows 2.
We write this as (T,2).         
(x) The coin shows tails. The die shows 3.
We write this as (T,3).         
(xi) The coin shows tails. The die shows 4.
We write this as (T,4).         
(xii) The coin shows tails. The die shows 5.
We write this as (T,5).         
(xiii) The coin shows tails. The die shows 6.
We write this as (T,6).
3. So the sample space can be written as:
S = {(H,B1), (H,B2), (H,B3), (H,W1), (H,W2), (H,W3), (H,W4), (T,1), (T,2), (T,3), (T,4), (T,5), (T,6)}
• There are 13 elements in S.

Solved Example 16.5
Consider the experiment in which a coin is tossed repeatedly until a head comes up. Write the sample space.
Solution:
1. Let us write the outcomes:
(i) Head is obtained in the first toss.
We write this as H.
(ii) Head is obtained only in the second toss.
We write this as TH.
(iii) Head is obtained only in the third toss.
We write this as TTH.
(iv) Head is obtained only in the fourth toss.
We write this as TTTH.
so on . . .
2. So the sample space can be written as:
S = {H, TH, TTH, TTTH, . . . }


Link to a few more solved examples is given below:

Exercise 16.1


In the next section, we will see Event.

Previous

Contents

Next

Copyright©2023 Higher secondary mathematics.blogspot.com